# AACR-Bench: Evaluating Automatic Code Review with Holistic Repository-Level Context

### v1

@@ Abstract
High-quality evaluation benchmarks are pivotal for deploying Large Language Models (LLMs) in Automated Code Review (ACR). However, existing benchmarks suffer from two critical limitations: first, the lack of multi-language support in repository-level contexts, which restricts the generalizability of evaluation results; second, the reliance on noisy, incomplete ground truth derived from raw Pull Request (PR) comments, which constrains the scope of issue detection. To address these challenges, we introduce AACR-Bench, a comprehensive benchmark that provides full cross-file context across multiple programming languages. Unlike traditional datasets, AACR-Bench employs an “AI-assisted, Expert-verified” annotation pipeline to uncover latent defects often overlooked in original PRs, resulting in a 285% increase in defect coverage. Extensive evaluations of mainstream LLMs on AACR-Bench reveal that previous assessments may have either misjudged or only partially captured model capabilities due to data limitations. Our work establishes a more rigorous standard for ACR evaluation and offers new insights on LLM based ACR, i.e., the granularity/level of context and the choice of retrieval methods significantly impact ACR performance, and this influence varies depending on the LLM, programming language, and the LLM usage paradigm e.g., whether an Agent architecture is employed.
@@ 1 Introduction
Nowadays, empowered by Large Language Models (LLMs), automated code review(ACR) technologies have been widely studied and adopted (Hou et al., 2024; Zhang et al., 2023). Leveraging the impressive comprehension capabilities of LLMs, researchers have explored various methods for the end-to-end generation of review comments, for example, the pre-training model schemes proposed by Tufano et al., fine-tuning strategies that balance cost and performance (Lu et al., 2023), and optimization mechanisms based on reinforcement learning (Le et al., 2022). Recently, Agent-based code review systems have also begun to emerge (Guo et al., 2025b; Ren et al., 2025).
@@
The flourishing research and adoption in ACR has intensified the need for robust benchmarks capable of comprehensively evaluating their performance. However, current evaluation frameworks still suffer from critical limitations, which can be summarized into two main issues.
@@
  • Incomplete Defect Annotation. In general, existing benchmarks directly use raw review comments from real-world historical PRs as the Ground Truth, resulting in inherent limitations regarding defect coverage within the dataset (Liu et al., 2025; Khoshnoud et al., 2022). This, in turn, makes it impossible to authentically and comprehensively and faithfully characterize a model’s ability to uncover potential defects in a typical code review scenarios.
@@
  • Restricted Context Scope. Many code defects manifest through cross-file dependencies, requiring systems to possess repository-level context awareness. While a few benchmarks provide such context (Zeng et al., 2025; Guo et al., 2025a), they are typically restricted to a single programming language (e.g., Python). This language homogeneity not only limits the generalizability of evaluation results but also introduces structural biases specific to that language, failing to reflect the multilingual nature of modern software projects.
@@
These deficiencies collectively undermine the reliability and scope of current ACR evaluations, highlighting the necessity for a more comprehensive and realistic benchmark.
@@
To address the aforementioned challenges, we propose AACR-Bench—a multilingual ACR benchmark that supports Repository-level Context-awareness. First, to address the issues of context and language diversity, AACR-Bench provides complete repository-level dependency information while broadly covering 10 mainstream programming languages. Secondly, to tackle the insufficiency of “Labeling” in existing datasets, we have constructed a high-quality hybrid dataset. In addition to incorporating $391$ real-world review comments extracted from raw GitHub PRs, we introduced high-density expert annotation data. Specifically, we enlisted over $80$ senior software engineers(each with over two years of company experience) to meticulously review and annotate $1505$ comments generated by two ACR systems powered by 6 different LLMs. This data augmentation strategy significantly enhances the Completeness of defect coverage, enabling AACR-Bench to quantify a model’s ability to mine potential defects more authentically and precisely than traditional datasets.
@@
  • We conduct a comprehensive empirical evaluation of the mainstream LLMs with AACR-Bench and reveal new and insightful observations, i.e., the granularity/level of context and the choice of retrieval methods significantly impact ACR performance, and this influence varies depending on the LLM, programming language, and the LLM usage paradigm (e.g., whether an Agent architecture is employed).
@@ 2 Related Work
We position this study by examining two key dimensions: the current state of large language model (LLM)-enabled automated code review, and the evolving landscape of corresponding specialized benchmark methodologies.
@@ 3.1 Overview of AACR-Bench
Figure 1: Overview of AACR-Bench
AACR-Bench is a repository-level benchmark tailored for evaluating the end-to-end generation performance of ACR methods/systems. Figure 1 show the overview of AACR-Bench. In general, AACR-Bench contains $200$ PRs and $1,505$ fine-grained review comments extracted and curated from $50$ popular repositories, covering $10$ mainstream programming languages. Besides, AACR-Bench incorporates a hybrid of model-augmented human reviews and fully model-generated reviews, all of which have been rigorously verified through human expert annotation to ensure benchmark validity.
@@
To align with the diff-oriented paradigm of modern code review, AACR-Bench designates each PR as an evaluation unit. During execution, the method under test traverses all Diff Hunks within a PR to generate review comments. Performance is evaluated by matching these generated comments against the Ground Truth (comprising 1,505 items) and calculating specific hit metrics.
@@ 3.2 Dataset Curation
Figure 2: Distribution of Review Comments in AACR-Bench . TS, JS, and Py stand for TypeScript, JavaScript, and Python, respectively. “Aug” denotes comments augmented from original PR reviews, while “Gen” denotes comments generated by LLMs.
As discussed in Section 2, existing benchmarks suffer from single-language confinement and incomplete ground truth derived solely from historical data. To bridge these gaps, we constructed a multilingual, repository-level benchmark with process depicted in Figure 4. Core considerations and treatments are listed as below:
@@
To ensure both the currency and diversity of our evaluation benchmark, we selected the top 10 programming languages based on the StackOverflow Developer Survey 2025, namely: JavaScript, Python, TypeScript, Java, C#, C++, C, PHP, Go, and Rust. For each language, we identified candidate repositories on GitHub that ranked within the top 2,000 in terms of both newly acquired stars and closed pull requests (PRs) from December 1, 2024, to December 1, 2025. From this candidate pool, we selected the top 5 repositories with the highest number of new stars. Consequently, our final dataset comprises 50 repositories across these 10 distinct languages (5 per language), serving as the source for extracting PRs and review comments.
PR Filtering and Review Comment Augmentation
@@
We extracted Pull Requests (PRs) from 50 selected repositories, discarding comments that did not identify substantive issues. To ensure data quality, we applied five filtering criteria: (1) PR titles and descriptions must be in English; (2) changed lines of code must be $\leq 1,000$ (aligning with Google’s code review practice for effective review); (3) the primary programming language of modified files must match the repository’s main language; (4) a PR must contain >2>2 inline comments, including at least one adopted comment that led to code modification; (5) changes detached from the project’s business context or lacking semantic meaning were excluded. Finally, to ensure the representativeness and diversity of the benchmark, we applied stratified sampling to the filtered candidates based on the repository, PR problem domain (Jimenez et al., 2023; Guo et al., 2025a), and change size, constructing a core dataset of 200 PRs. As code review often involves multi-turn dialogue, directly using raw comments may lose context or introduce noise. Focusing on the revision with the most inline comments per PR, we employed an LLM to perform deep semantic analysis on review threads. This allowed us to extract confirmed code defects from multi-turn interactions and reorganize them into “Augmented Review Comments.”
@@
Review Completion and Expert Annotation To address the Label Incompleteness stemming from the often under-reviewed nature of GitHub PRs (Bacchelli & Bird, 2013), we leveraged LLMs to comprehensively supplement the dataset. We constructed a generation matrix comprising seven mainstream open- and closed-source models() to mitigate single-model bias and ensure output diversity. Review comments were generated in parallel through two heterogeneous frameworks: an internal review system and the open-source agent, Claude Code. Following semantic de-duplication, the generated comments were merged with the augmented human reviews to form a candidate set for verification. We then enlisted 80 senior software engineers (each with over two years of experience) to perform rigorous Human Annotation. Annotators validated the correctness of the comments and categorized the issue types. Distinguishing our work from existing benchmarks, we innovatively annotated the context level required to formulate each comment, enabling the assessment of detection difficulty across varying context dependencies. This step yielded a total of 1,505 review comments, including 391 augmented from original reviews and 1,114 generated by LLMs, representing a 285% increase in problem coverage. Figure 2 illustrates the language distribution and the proportion of augmented human versus model-generated comments.
@@ 3.3 Comparisions with Existing Datasets
Dataset Multi-Languages support Repo-level context support Source of review comments Context level annotation
@@
A brief comparison between existing ACR benchmarks and ours are shown in Table 2. The primary distinction between our benchmark and existing works is that ours is the first to provide repository-level context within a multi-language environment. While Codefuse-CR-Bench (Guo et al., 2025a) and SWR-Bench (Zeng et al., 2025) provide repository-level context (including PR information and full repository code) to assess tools under complex, realistic conditions, they are exclusively limited to Python repositories. This confinement restricts the generalizability of their results and risks introducing structural bias inherent to a single language. Conversely, although ContextCRBench (Hu et al., 2025) spans multiple programming languages, it provides only class- or function-level context(equals to file level in our study), thereby limiting its ability to evaluate tools in complex scenarios. Similarly, CodeReviewer—currently the most prevalent benchmark—covers multiple languages but relies solely on the code Diff as context. In contrast, our work synthesizes the strengths of these predecessors. We evaluate automated code review tools using rich, repository-level context across 10 popular programming languages (JavaScript, Python, TypeScript, Java, C#, C++, C, PHP, Go, and Rust). This context includes the ‘PR Title’, ‘Description’, and the complete code repository.
A second critical distinction lies in the origin and quality of the Ground Truth review comments. Existing datasets such as Codefuse-CR-Bench, SWR-Bench, and CodeReviewer derive their ground truth directly from raw review comments found in GitHub PRs. However, several studies revealed that the issues in these PRs are often insufficiently identified (Bacchelli & Bird, 2013), leading to potentially bias benchmark results. Furthermore, raw data is inherently noisy; comments frequently consist of redundant discussions, informational synchronization, or non-substantive approval messages (e.g., ‘LGTM’, ‘Merge’) (Rong et al., 2024). In contrast, we employ a more rigorous construction methodology. First, we use advanced Large Language Models (LLMs) to refine raw comments, extracting specific technical issues to eliminate conversational noise. Second, to maximize coverage of potential defects, we expand the ground truth using two distinct review systems powered by $6$ advanced open-source and proprietary LLMs(Qwen3-Coder-480B-A35B-Instruct, GLM-4.7, DeepSeek-V3.2, Claude-4.5-Sonnet, GPT-5.2, Gemini-3-Pro). Crucially, all review comments—whether refined from human inputs or newly generated by models—are meticulously annotated by professional software engineers to ensure their reliability.
@@ 4.1 Evaluation Settings
Models We selected the latest models from mainstream open-source providers and the most recent versions of major commercial large models. As the result, our experiment comprises three open-source models (Qwen3-Coder-480B-A35B-Instruct, DeepSeek-V3.2, GLM-4.7) and two commercial models (GPT-5.2, Claude-4.5-Sonet). We evaluated all selected models using our complete dataset. For simplicity, Qwen3-Coder-480B-A35B-Instruct will be called Qwen-480B-Coder in the following contents.
Context Retrival Method One distinctive feature of our dataset is the provision of repository-level contextual dependencies. However, different retrieval methods also significantly influence the model’s reasoning performance (Zhang et al., ; Liu et al., 2023b). In this experiment, the following approaches were adopted:
@@
  • BM25: A classic text similarity method (Robertson et al., 2009), also applied in a previous related study (Guo et al., 2025a).
@@
Table 3: Performance Comparison (%) of Models Across Different Types
Type Model Avg Comments Recall (%) Precision (%) F1 (%)
@@ 4.2 Main Results
This section presents a comprehensive evaluation of diverse Code Review methodologies and models on our benchmark. We specifically analyze the impact of context retrieval, model architectures, and agent frameworks on the overall quality of generated reviews. It is important to note that both the current practical applications of ACR (Sadowski et al., 2015; Distefano et al., 2019) and prior research (Tao et al., 2012) consistently highlight the critical role of contextual dependency information in ACR. Therefore, all experiments in this paper are based on the assumption that this represents the geneunin paradigm established for ACR. In short, unless serving as a comparative basis, we do not evaluate the performance of LLMs on ACR tasks without contextual information.
@@ 4.2.1 ACR performance benchmarked with a more sufficient exposure to defects
Agent methods (e.g., Claude Code) produce far fewer review comments per patch ($0.08\sim 0.15$) than traditional approaches. For instance, Claude-4.5-Sonnet in Agent mode achieves a Precision of $39.90\%$, drastically higher than its $8.70\%$ in “No context” mode, but at the cost of much lower Recall ($10.10\%$). This suggests that while Agents excel in precision-sensitive settings, they may miss many potential issues—a tendency we hypothesize stems from “contextual tunnel vision” (Liu et al., 2023a) induced by their targeted retrieval mechanism. In contrast, traditional methods often generate excessive comments (e.g., GPT-5.2 averages $2.47$ per patch), which can bury valuable insights in noise.
Further analysis reveals that the effectiveness of Agent-based methods exhibits a pronounced dependency on model-specific characteristics, a sensitivity not observed in other context retrieval strategies. While Claude-4.5-Sonnet achieved a remarkable $39.90\%$ Precision in Agent mode, other models failed to replicate this performance leap upon switching to Agent mode, instead suffering from varying degrees of performance collapse. Taking GPT-5.2 for example, despite demonstrating robust baseline performance in “no context” mode , its performance plummeted in Agent mode, with Precision dropping to $9.90\%$ and Recall falling to a mere $2.99\%$. This stark contrast indicates that general-purpose capabilities may not straightforwardly translate into competence for Agent-based ACR scenarios (Liu et al., 2023c).
@@
Table 3 further indicates that context retrieval is not universally beneficial. For models with strong inherent reasoning, naive RAG can introduce harmful noise, while various open-source models show distinct preferences for specific retrieval strategies. For instance, Claude-4.5-Sonnet performs robustly in “o context” mode (F1=$14.46$). However, providing top-3 contexts via BM25 causes performance to drop sharply (F1=$9.98$, a $31\%$ decrease; Precision falls from $8.70\%$ to $5.80\%$). Embedding-based retrieval similarly degrades its results. Different models favor different strategies. DeepSeek-V3.2 achieves its best performance under BM25: Avg. Comments per patch decreases from 2.29 (“No context”) to $0.89$, while Precision jumps from $5.10\%$ to $10.90\%$, yielding an F1 of $15.59$. Conversely, Qwen-480B-Coder peaks using Embedding-based retrieval (F1=$14.36$), significantly outperforming its BM25 score ($11.69$). Thus, no single retrieval method is optimal for all models; performance is highly dependent on the specific combination of model and retrieval mode.
@@
The Agent-based approach generates significantly fewer review comments compared to other methods. Its effectiveness is highly dependent on model characteristics, with performance varying significantly across different models when using this approach. Furthermore, different models respond differently to context retrieval methods, e.g., pairing Claude-4.5-Sonnet with Agent frameworks, DeepSeek with BM25, and Qwen with Embedding consistently present optimal performance.
Table 4: ACR performance with Different Contextual levels across Methods (Recall only)

### v2

@@ Abstract
High-quality evaluation benchmarks are pivotal for deploying Large Language Models (LLMs) in a certain applicaiton field. However, existing benchmarks on Automated Code Review (ACR) suffer from two critical limitations: first, the reliance on noisy, incomplete ground truth derived from raw Pull Request (PR) comments, which constrains the scope of issue detection; second, the lack of multi-language support in repository-level contexts, which restricts the generalizability of evaluation results. To address these challenges, we introduce AACR-Bench, a comprehensive benchmark that provides full cross-file context across multiple programming languages. Unlike traditional datasets, AACR-Bench employs an “AI-assisted, Human expert-verified” annotation pipeline to uncover latent defects often overlooked in original PRs, resulting in a 285% increase in issue coverage. Extensive evaluations of mainstream LLMs on AACR-Bench reveal that previous assessments may have either misjudged or only partially captured model capabilities due to data limitations. Our work establishes a more rigorous standard for ACR evaluation and offers new insights on LLM based ACR, i.e., the granularity/level of context and the choice of retrieval methods significantly impact ACR performance, and this influence varies depending on the LLM, programming language, and the LLM usage paradigm e.g., whether an Agent architecture is employed. The code, data, and other artifacts of our evaluation set are available at Github [^1] https://github.com/alibaba/aacr-bench.
@@ 1 Introduction
Nowadays, empowered by Large Language Models (LLMs), automated code review(ACR) technologies have been widely studied and adopted (Hou et al., 2024; Zhang et al., 2023). Leveraging the impressive comprehension capabilities of LLMs, researchers have explored various methods for the end-to-end generation of review comments, for example, the pre-training model schemes proposed by Tufano et al., the fine-tuning strategies that balance cost and performance (Lu et al., 2023), and the optimization mechanisms based on reinforcement learning (Le et al., 2022). Recently, Agent-based code review systems have also begun to emerge (Guo et al., 2025b; Ren et al., 2025).
@@
The flourishing research and adoption in ACR has intensified the need for credible benchmarks capable of comprehensively evaluating their performance. However, current evaluation frameworks still suffer from critical limitations, which can be summarized into two main issues.
@@
  • Incomplete Issue Annotation. In general, existing benchmarks directly use raw review comments from real-world historical PRs as the Ground Truth, resulting in inherent limitations regarding issue coverage within the dataset (Liu et al., 2025; Khoshnoud et al., 2022). This, in turn, makes it impossible to authentically and faithfully characterize a model’s ability to uncover potential issues in a real-world code review scenarios.
@@
  • Restricted Context Scope. Many code defects are inherently cross-file, requiring review systems to access repository-level context for accurate detection. Although some existing benchmarks do offer cross-file context awareness (Zeng et al., 2025; Guo et al., 2025a), they are predominantly limited to a single programming language, such as Python. This language-specific focus not only constrains the generalizability of evaluation findings but may also introduce structural biases tied to the linguistic characteristics of that language. Consequently, such benchmarks fail to adequately represent the multilingual reality of contemporary software development.
@@
These deficiencies collectively undermine the credibility and scope of current ACR evaluations, highlighting the necessity for a more comprehensive and crediable benchmark.
@@
To address the aforementioned challenges, we propose AACR-Bench—a multilingual ACR benchmark that supports Repository-level Context-awareness. First, to address the limited issue annotation in existing datasets, we constructed a high-quality hybrid corpus. Beyond collecting 391 real-world review comments from GitHub PRs, we incorporated extensive manual issue annotation: 80 senior software engineers (each with 2+ years of industry experience) meticulously reviewed 2,145 comments generated by two ACR systems across six LLMs. This annotation strategy substantially improves issue coverage and enables AACR-Bench to more accurately and comprehensively assess a model’s capability to uncover latent issues compared to conventional datasets. Secondly, to address the issues of context and language diversity, AACR-Bench provides complete repository-level dependency information while broadly covering 10 mainstream programming languages. The full AACR-Bench dataset can be retrieved in the supplementary materials of this paper and will be open-sourced later on.
@@
  • We combine multi-model generation and large-scale human annotation to improve issue coverage so as to establish a better groundtruth dataset.
@@
  • We conduct a comprehensive empirical evaluation of the mainstream LLMs with AACR-Bench and reveal new insights towards mainstream LLM’s ACR capabilities.
@@ 2 Related Work
We position this study by examining two key dimensions: the current state of large language model (LLM)-enabled ACR, and the evolving landscape of corresponding specialized benchmark methodologies.
@@ 3.1 Overview of AACR-Bench
Figure 1: Overview of AACR-Bench. AACR-Bench contains $200$ PRs and $1,505$ fine-grained review comments extracted and curated from $50$ popular repositories, covering $10$ mainstream programming languages.
AACR-Bench is a repository-level benchmark tailored for evaluating the end-to-end performance of ACR methods/systems. Figure 1 show the overview of AACR-Bench. In general, AACR-Bench contains $200$ PRs and $1,505$ fine-grained review comments extracted and curated from $50$ popular repositories, covering $10$ mainstream programming languages. Besides, AACR-Bench incorporates a hybrid of model-augmented human reviews and fully model-generated reviews, all of which have been rigorously verified through human expert annotation to ensure benchmark credibility.
@@
To align with the diff-oriented paradigm of modern code review, AACR-Bench designates each PR as an evaluation unit. During execution, the ACR method under assessment traverses all diff Hunks within a PR to generate review comments. Performance is evaluated by matching these generated comments against the Ground Truth (comprising 1,505 items) and calculating specific accuracy metrics(e.g., Precision, Recall and F1-score.)
@@ 3.2 Dataset Curation
Figure 2: Distribution of Review Comments in AACR-Bench . TS, JS, and Py stand for TypeScript, JavaScript, and Python, respectively. “Aug” denotes comments augmented from original PR reviews, while “Gen” denotes comments generated by the 6 LLMs.
As discussed in Section 1, existing benchmarks suffer from single-language confinement and incomplete ground truth derived solely from historical PR data. To bridge these gaps, we constructed AACR-Bench  a multilingual, repository-level benchmark. Core considerations and treatments are listed as below:
@@
To ensure both the currency and diversity of our evaluation benchmark, we selected the top 10 programming languages based on the StackOverflow Developer Survey 2025222https://survey.stackoverflow.co/2025/technologyÅ, namely: JavaScript, Python, TypeScript, Java, C#, C++, C, PHP, Go, and Rust. For each language, we identified candidate repositories on GitHub that ranked within the top 2,000 in terms of both newly acquired stars and closed PRs from December 1, 2024, to December 1, 2025. From this candidate pool, we selected the top 5 repositories with the highest number of new stars. Consequently, our final dataset comprises 50 repositories across these 10 distinct languages (5 repositories per language), serving as the original source for extracting PRs and review comments.
PR filtering and comment augmentation
@@
With the extracted PRs from 50 selected repositories, we discard comments that did not identify explicit issues. To ensure data quality, we applied five filtering criteria: (1) PR titles and descriptions must be in English; (2) changed lines of code must be $\leq 1,000$ (aligning with Google’s code review practice for effective review (Solmaz, 2025)); (3) the primary programming language of modified files must match the repository’s main language; (4) a PR must contain >2>2 inline comments, including at least one adopted comment that led to code modification; (5) changes detached from the project’s business context or lacking semantic meaning were excluded. Finally, to ensure the representativeness and diversity of the benchmark, we applied stratified sampling to the filtered candidates based on the repository, PR problem domain (Jimenez et al., 2023; Guo et al., 2025a), and change size, constructing a core dataset of 200 PRs. As code review often involves multi-turn dialogue, directly using raw comments may lose context or introduce noise. Focusing on the revision with the most inline comments per PR, we employed an LLM to perform deep semantic analysis on review threads. This allowed us to extract confirmed code defects from multi-turn interactions and reorganize them into “Augmented Review Comments.”
Review completion and expert annotation
@@
To address the Label Incompleteness stemming from the often under-reviewed nature of GitHub PRs (Bacchelli & Bird, 2013), we leveraged LLMs to comprehensively supplement the review comments for each PR. We constructed a generation matrix comprising 6 mainstream open- and closed-source models(Claude-4.5-Sonnet, Qwen3-Coder-480B-A35B-Instruct (Team, 2025), GPT-5.2, Deepseek-V3.2 (DeepSeek-AI, 2025), GLM-4.7 (Team et al., 2025), Gemini-3-Pro) to mitigate single-model bias and ensure output diversity. Review comments were generated in parallel through two heterogeneous frameworks: an internal review system and the open-source agent, Claude Code. Following semantic de-duplication, the generated comments were merged with the augmented human reviews to form a candidate set for human verification later on. The collection was subjected to rigorous human annotation by 80 senior software engineers, each with over two years of professional experience. Every comment was annotated by at least two annotators independently, with any discrepancies resolved through discussion by a six-member core team. Annotators validated the correctness of the comments and categorized the issue types(Sun et al., 2025). Distinguishing our work from existing benchmarks, we innovatively annotated the context level required to formulate each comment, enabling the assessment of detection difficulty across varying context dependencies. This step yielded a total of 1,505 review comments, including 391 augmented from original reviews and 1,114 augmented by LLMs and human experts, representing a 285% increase in issue coverage. Figure 2 illustrates the language distribution and the proportion of augmented human versus model-generated comments.
@@ 3.3 Comparisions with Existing Datasets
Dataset Multi-Languages support Repo-level context support Source of review comments Context scope annotation
@@
A brief comparison between existing ACR benchmarks and ours are shown in Table 2. Our benchmark distinguishes itself as the first to provide repository-level context within a multi-language environment, synthesizing the strengths of existing works. While Codefuse-CR-Bench (Guo et al., 2025a) and SWR-Bench (Zeng et al., 2025) offer repository contexts, they are restricted to Python; conversely, multi-language datasets like ContextCRBench (Hu et al., 2025) and CodeReview (Li et al., 2022) are limited to file-, function- or even diff-level contexts. We bridge this gap by evaluating tools across 10 popular programming languages with comprehensive context, including PR metadata and complete code repositories.
Furthermore, we address the noise and incompleteness inherent in raw GitHub data used by prior benchmarks (Guo et al., 2025a; Zeng et al., 2025; Hu et al., 2025; Li et al., 2022). Instead of relying on raw extraction, we employ a rigorous construction methodology: we use LLMs to augment technical issues from human inputs and expand defect coverage using six advanced models (including Qwen3-Coder-480B, GPT-5.2, and Gemini-3-Pro). To ensure reliability, all comments—whether augmented or generated—are meticulously verified by a large-scale team consisting of 80 professional software engineers.
Finally, beyond standard PR and comment labels (Guo et al., 2025a; Zeng et al., 2025; Hu et al., 2025), we innovatively annotate the required context scope for each review comment. This novel dimension enables a granular analysis of an ACR method’s capacity to detect issues dependent on varying context depths and scope.
@@ 4.1 Evaluation Settings
Models We selected the latest models from mainstream open-source providers and the most recent versions of major commercial large models. As the result, our experiment comprises three open-source models (Qwen3-Coder-480B-A35B-Instruct (Team, 2025), DeepSeek-V3.2 (DeepSeek-AI, 2025), GLM-4.7 (Team et al., 2025)) and two commercial models (GPT-5.2, Claude-4.5-Sonet). We evaluated all selected models using our complete dataset. For simplicity, Qwen3-Coder-480B-A35B-Instruct will be called Qwen-480B-Coder in the following contents.
Context Retrival Method One distinctive feature of our dataset is the provision of repository-level contextual dependencies. However, different retrieval methods also significantly influence the model’s reasoning performance (Zhang et al., ; Liu et al., 2023b). In this experiment, the following methods were adopted:
@@
  • BM25: A classic text similarity method (Robertson et al., 2009), also applied in a previous similar study (Guo et al., 2025a).
@@
Table 3: Performance Comparison (%) of Models Across Different ACR Methods
Method Model Avg Comments Recall (%) Precision (%) F1 (%)
@@ 4.2 Main Results
This section presents a comprehensive evaluation of diverse ACR methodologies and models on our benchmark. We specifically analyze the impact of context retrieval, model architectures, and agent frameworks on the overall quality of generated reviews. It is important to note that both the current practical applications of ACR (Sadowski et al., 2015; Distefano et al., 2019) and prior research (Tao et al., 2012) consistently highlight the critical role of contextual dependency information in ACR. Therefore, all experiments in this paper are based on the assumption that this represents the geneunin paradigm established for ACR. In short, unless serving as a comparative basis, we do not evaluate the performance of LLMs on ACR tasks without contextual information.
@@ 4.2.1 ACR performance benchmarked with a more sufficient exposure to defects
Agent-based methods (e.g., Claude Code) consistently generate substantially fewer review comments per patch ($0.08\sim 0.15$) compared to traditional approaches. For instance, Claude-4.5-Sonnet in Agent mode attains a Precision of $39.90\%$, far exceeding its $8.70\%$ in “No context” mode, yet accompanied by a notably lower Recall of $10.10\%$. This indicates that while Agents excel in precision-critical scenarios, they may overlook many potential defects—a tendency attributable to the “contextual tunnel vision” effect induced by their focused retrieval mechanism (Liu et al., 2023a). In contrast, traditional methods tend to produce an excess of comments (e.g., GPT-5.2 averages $2.47$ per patch), which can obscure meaningful insights amidst noise. Furthermore, the effectiveness of Agent-based methods exhibits strong model-specific dependency, a sensitivity not observed with other retrieval strategies. Although Claude-4.5-Sonnet achieves exceptional precision in Agent mode, other models do not show similar gains and may even experience performance degradation. For example, GPT-5.2, despite strong baseline performance in “no context” mode, suffers a sharp decline in Agent mode, with Precision falling to $9.90\%$ and Recall dropping to $2.99\%$. This contrast suggests that general-purpose model capabilities do not directly translate to proficiency in Agent-based ACR tasks (Liu et al., 2023c).
@@
Table 3 further indicates that context retrieval is not universally beneficial. For models with strong inherent reasoning, naive RAG can introduce harmful noise, while various open-source models show distinct preferences for specific retrieval strategies. For instance, Claude-4.5-Sonnet performs robustly in “no context” mode (F1=$14.46$). However, providing top-3 contexts via BM25 causes performance to drop sharply (F1=$9.98$, a $31\%$ decrease; Precision falls from $8.70\%$ to $5.80\%$). Embedding-based retrieval similarly degrades its results. Different models favor different strategies. DeepSeek-V3.2 achieves its best performance under BM25: Avg. Comments per patch decreases from 2.29 (“No context”) to $0.89$, while Precision jumps from $5.10\%$ to $10.90\%$, yielding an F1 of $15.59$. Conversely, Qwen-480B-Coder peaks using Embedding-based retrieval (F1=$14.36$), significantly outperforming its BM25 score ($11.69$). Thus, no single retrieval method is optimal for all models; performance is highly dependent on the specific combination of model and retrieval mode.
@@
The Agent-based approach consistently produces far fewer review comments than other methods, and its effectiveness is strongly model-dependent. Performance varies markedly across different LLMs under this paradigm. Furthermore, different models respond differently to context retrieval methods, e.g., pairing Claude-4.5-Sonnet with Agent frameworks, DeepSeek with BM25, and Qwen with Embedding consistently present optimal performance.
Table 4: Performance on Finding Issues Using Different Contextual Scopes Across Various ACR Methods (Recall only)